$SPAD/src/input d01asf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
e:Expression Float:=1/sqrt(X)
f:ASP1(G):=retract e
a:SF:=1.0e-12
omega:SF:=%pi/2
key:=1
epsabs:SF:=1.0e-3
limlst:=50
liw:=400
lw:=2*liw
result:=d01asf(a,omega,key,epsabs,limlst,lw,liw,-1,f)

———-

References^{[1] nothing}

4